Suppose the risk-free rate is 3.21% and an analyst assumes a market risk premium of 7.86%. Firm A just paid a dividend of $1.16 per share. The analyst estimates the β of Firm A to be 1.33 and estimates the dividend growth rate to be 4.63% forever. Firm A has 288.00 million shares outstanding. Firm B just paid a dividend of $1.68 per share. The analyst estimates the β of Firm B to be 0.79 and believes that dividends will grow at 2.79% forever. Firm B has 195.00 million shares outstanding. What is the value of Firm B?
Answer #1

As per CAPM, required rate of return = risk free rate+ beta*market risk premium

Firm B = 3.21% + 0.79*7.86% = 9.4194%

Value per share = expected dividend/(required return - growth rate)

= 1.68(1.0279)/(9.4194%-2.79%)

=$26.05

Value of firm = value per share*number of shares

=26.05*195 million

=$5,079.75 million